Circular No. 8003

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 T7 (LINEAR) Another object reported as asteroidal by the LINEAR survey has been

found to be cometary in appearance. P. Birtwhistle, Great Shefford, U.K., writes that CCD images taken on Oct. 28.0 UT show the object to be slightly ‘softer’ than stars of similar brightness. T. B. Spahr reports that Mt. Hopkins images taken on Oct. 29.4 show the object to be very slightly diffuse, again with it being slightly larger than the FWHM of nearby stars.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Oct. 14.41856  6^{h}03^{m}49.^{s}91  +27^{o}50′02′′.7  17.5

The available astrometry (some giving the comet’s magnitude as bright as 16.5), including an Oct. 12 prediscovery observation, is given on MPEC 2002-U43, together with the preliminary parabolic orbital elements below. It is possible that this comet could reach naked-eye brightness around April– June 2004.

T = 2004 Apr. 23.685 TT  ω = 157.^{o}770
Ω = 94.865 }2000.0
q = 0.61523 AU  i = 160.594
2002TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 23 6^{h}00.^{m}74 +28^{o}00.^{′}3 6.273 6.813 119.^{o}1 7.^{o}3 16.3
Nov. 2 5 55.83 +28 12.1 6.041 6.724 130.1 6.5 16.2
12 5 49.50 +28 22.9 5.831 6.634 141.4 5.3 16.0
22 5 41.86 +28 31.5 5.648 6.544 153.0 3.9 15.9
Dec. 2 5 33.10 +28 37.0 5.497 6.453 164.6 2.3 15.8
12 5 23.53 +28 38.5 5.381 6.361 174.1 0.9 15.7
22 5 13.57 +28 35.4 5.303 6.269 168.4 1.8 15.6

V4744 SAGITTARII = POSSIBLE NOVA IN SAGITTARIUS N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4744 Sgr has been given to this object (cf. IAUC 8001).

SUPERNOVAE 2002hc AND 2002hd W. Li provides revised positions (equinox 2000.0) for the new LOTOSS

supernovae reported on IAUC 7999: SN 2002hc, α = 8^{h}17^{m}06^{s}.15, δ = −27^{o}27′27′′.8. SN 2002hd, α = 8^{h}54^{m}03^{s}.39, δ = −7^{o}11′21′′.5.

2002 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green